OpenStudy (anonymous):

What is the radius for the circle given by the equation x^2+(y-1)^2=12? do i just find the square root of the 12, or is there another step i need to take.

OpenStudy (raden):

yeah, u are right.. r^2 = 12, so r=sqrt(12) if u want simplify it can be : r=sqrt(12) = sqrt(4*3) = sqrt(4) * sqrt(3) = 2sqrt(3)